# Synchronization of Dirac-Bianconi driven oscillators

### v1

@@ I Introduction
In complex systems, the dynamics is shaped by the interactions among the elementary units and such interactions have been mostly assumed to be pairwise (two-body). However, in recent years, increasing evidence emerged towards the fundamental role that higher-order interactions play in describing real-world systems battiston2020networks ; bianconi2021higher ; bick2023higher ; boccaletti2023structure . This triggered the interest of scholars in nonlinear dynamics, particularly due to the interplay between dynamics and higher-order topology natphys ; muolo2024turing ; millan2025topology . In fact, it has been shown that higher-order interactions enrich the dynamics of the systems, with applications regarding synchronization tanaka2011multistable ; skardal2020higher ; millan2020explosive ; lucas2020multiorder ; wang2021collective ; skardal2021higher ; gambuzza2021stability ; gallo2022synchronization , random walks schaub2020random ; carletti2020random , pattern formation carletti2020dynamical ; muologallo , chimera states kundu2022higher ; muolo2024phase , opinion dynamics iacopini2019simplicial ; deville2020consensus ; ziegler2022balanced ; schawe2022higher , and control de2022pinning ; della2023emergence ; muolo2024pinning , to name a few. In all of the above works, the authors showed that certain dynamics are only made possible by the higher-order interaction topology, namely, hypergraphs and simplicial complexes, which are extensions of networks. This is particularly interesting, as the complexity of the system lies not in the model, but rather in the interactions between the elementary units, and even very simple settings yield rich dynamics when higher-order interactions are present kuehn2021universal ; leon2024 .
@@
In the study of dynamics on networks, the classical approach relies on node-based dynamics, where to each node is associated a variable, e.g., the phase of an oscillator, or the state of a neuron. Higher-order interactions in the sense of many-body (group) interactions persist in such node-based paradigm: in fact, the dynamical variables remain anchored to the nodes, while what change is the interactions between them, mediated by hyperedges of higher orders, i.e., group interactions. However, there are systems which do not fit well in this paradigm, and some examples come from collaboration networks patania2017shape , brain networks faskowitz2022edges , climate science singh2020fingerprint , and transportation networks katifori2010damage . To describe and analyze such systems, scholars have developed a framework that goes beyond node-based dynamics, using tools from algebraic topology and exterior calculus bianconi2021higher , which allow us to consider the dynamics of topological signals, i.e., state variables defined not only on the nodes, but also on links, triangles, and higher-order simplices. The structure on which the dynamics is defined is called simplicial complex, which is a particular case of hypergraph with a closure relation, i.e., each $(d+1)$-dimensional simplices needs to be bounded by all $d$-simplices, and so on. As a simple example, a $2$-simplex (i.e., a triangle), needs to be bounded by all $1$-simplices (i.e., the links), which have to be bounded by all the 00-simplices (i.e., the nodes). In this notation, a network can be considered as a $1$-simplicial complex, i.e., a simplicial complex made only of $1$- and 00-simplices. Moreover, though the Dirac-Bianconi operator bianconi2021topological , variables on different simplices of $1$ dimension up or down can be coupled. In this new framework, several classical nonlinear dynamics on networks model, such as the Kuramoto model kuramoto1975 , Benjamin-Feir instability nakao2014complex , the Master Stability Function fujisaka1983stability ; pecora1998master , and Turing pattern formation NM2010 ; muolo2024turing , have been extended for topological signals on simplicial complexes, namely, the Topological Kuramoto model  millan2020explosive ; Calmon2021topological ; ghorbanchian2021higher ; li2023synchronization ; arnaudon2022connecting ; nurisso2024unified , Global Topological Synchronization carletti2023global ; wang2025higher ; carletti2025global , and Topological Turing patterns Giambagli2022diffusion ; muolo2024three . The study of patterns and oscillations of topological signals in higher-order simplices is particularly interesting for the study of edge-signals in the brain faskowitz2022edges .
@@
In this work, we introduce a setting that is complementary to the above works. Before that, let us take a step back to the node-based framework. When studying synchronization dynamics on networks, it is common to view each node as a self-sustained oscillator, i.e., with a stable limit cycle, and then consider the synchronization between them pikovsky2001synchronization . When considering such systems, there is a powerful technique, called phase reduction, which allows to reduce each high dimensional oscillators to a phase variable nakao16 . In general, the application of phase reduction yields simple phase models that are easier to understand and that have been widely studied. In fact, the Kuramoto model is derived through phase reduction Kuramoto_book and provides a simple and general mechanism for the emergence of synchronization Kuramoto_book ; acebron2005kuramoto ; rodrigues2016kuramoto . In recent years, extensions of this method have been considered, such as corrections via an expansion of the (small) coupling strength leon19 ; leon22a ; Mau23 ; bick2023higher , the dynamics of the amplitude variables under non-weak coupling kurebayashi2013phase ; Wilson_2016 ; shirasaka2017phase , or mean field variables leon20 , which allowed to overcome some of the limitations and extend its applicability. All the above references consider the nodes as oscillators. However, periodic dynamics can emerge also in more complex settings. For example, most neurons are not oscillatory, but oscillatory patterns are observed at larger scale wilson1972excitatory ; brunel1999fast . This can be modeled by a network of interacting excitable units, whose isolated dynamics are not oscillatory, but exhibit self-sustained oscillations as a result of the coupling kawamura2011collective ; pazo2016quasiperiodic ; devalle2017firing ; nakao2018phase ; politi2018collective ; di2018transition ; nakao2021sparse . In particular, authors of Ref. nakao2018phase derived a phase description of oscillatory networks and studied their synchronization through phase reduction theory.
@@
Inspired by the latter work, we further extend this framework by considering networks where the dynamical variables lie on the nodes and the links, and are solely coupled via the Dirac-Bianconi operator. When such system exhibit self-sustained oscillations, we call them Dirac-Bianconi driven oscillators, because their periodic behavior is induced by the Dirac-Bianconi operator. We then give a full description of such structures and study their synchronization through phase reduction theory.
@@
The paper is organized as follows. In the next Section we will give the theoretical background on topological signals and the Dirac-Bianconi operator. In Section III, we first introduce and characterize a Dirac-Bianconi driven oscillator, and then study numerically the synchronization of two coupled oscillators. Finally, in Sec. IV, we give a phase description of such systems that explains the numerical results.
@@ II Theory of topological signals and the Dirac-Bianconi operator
In this Section we present a basic introduction to simplicial complexes and topological signals, limiting ourselves to define what is needed to characterize Dirac-Bianconi driven oscillators. For a more detailed discussion that includes other important concepts, the interested reader may consult more specialized references bianconi2021higher ; millan2025topology ; Lim2020 ; schaub2021signal .
A $d$-dimensional simplex is a set of $d+1$ nodes that models a higher-order interaction among $(d+1)$ entities or represents a discrete object of dimension $n$. A 00-simplex is a node, a $1$-simplex is a link, a $2$-simplex is a triangle, and so on. The faces of a $d$-simplex are all its $(d-1)$-dimensional subsimplices. For instance, the faces of a triangle (2-simplex) are its three links (1-simplices), and each link has two node (0-simplex) faces. Except for the 00-simplices, each simplex has an arbitrary orientation, i.e., each simplex is labeled by an ordered list of nodes. For example, the (symmetric) link connecting nodes $i$ and $j$ can be labeled as $[i,j]$; then $[j,i]=[i,j]$. Linear combinations of $d$-simplices are also called $d$-chains. A simplicial complex is built by attaching simplices by their faces. Note that, in a simplicial complex, the existence of a $d$-simplex implies the existence of all its faces. Given $N_{k}$ $k$-simplices and $N_{k-1}$ $(k-1)$-simplices, the boundary operator $\mathbf{B}_{k}$ is a $N_{k-1}\times N_{k}$ matrix defined by:
$\mathbf{B}_{k}(i,j)=\begin{cases}1&\text{if the $i$-th $(k-1)$-simplex is a face of the $j$-th $k$-simplex with same orientation,}\\ -1&\text{if the $i$-th $(k-1)$-simplex is a face of the $j$-th $k$-simplex with opposite orientation,}\\ 0&\text{if the $i$-th $(k-1)$-simplex is NOT a face of the $j$-th $k$-simplex.}\end{cases}$ (1)
Note that the boundary operator $\mathbf{B}_{k}$ changes when the orientation is changed, but the structure of the simplicial complex does not change; one can understand this by thinking of Cartesian coordinates: the axis can be changed, but this does not change the structure of the space. Note also that, in the case of network, i.e., consisting only of 00-simplices and $1$-simplices, $\mathbf{B}_{1}$ is the well-known incidence matrix newmanbook .
We now define topological signals (also called cochains) as functions defined on simplices (chains). A $k$-dimensional signal is a function assigning a real value to each $k$-simplex $\sigma^{i}_{k}$. The values can be arranged in a vector [^1] Note that the dimension of the topological signal and the dimension of the simplex are not related. As one can define an $n$-dimensional vector on the nodes (00-simplices), the same can be done on any $k$-simplex. $\vec{x}\in\mathbb{R}^{N_{k}}$, where each entry is $x_{i}=x(\sigma^{i}_{k})$. These signals must be orientation-sensitive, i.e., $x(\sigma_{k})=-x(-\sigma_{k})$, to ensure the invariance under the change of the arbitrary orientation carletti2023global .
@@
The boundary operator (1) connects topological signals (cochains) of consecutive dimensions: $\mathbf{B}_{k}$ on a $k$-cochain $\vec{x}$, it gives a $(k-1)$-cochain $\vec{z}=\mathbf{B}_{k}\vec{x}$, while acting with the transpose gives a $(k+1)$-cochain $\vec{g}=\mathbf{B}_{k+1}^{\top}\vec{x}$. The transpose of the boundary is known as the coboundary operator. One cannot connect topological signals of dimensions $k$ and $k\pm 2$ because the (co)boundary of a (co)boundary is zero bianconi2021higher .
For sake of simplicity, let us consider the case of a $1-$simplicial complex, i.e., a network, hence, only $0-$simplices (nodes) and $1-$simplices (links) are involved, with their respective topological signals [^2] Note that the terms cochains, which we will mainly use from now on, and topological signals are equivalent., namely, $0-$cochains $\vec{u}$ (state variables on the nodes) and $1-$cochains $\vec{v}$ (state variables on the links). Note that the same formalism can be used to consider any dimension, as long as they are consecutive, given that Dirac-Bianconi operator connects topological signals of consecutive dimension. The state of the system is described by the topological spinor $\vec{w}$, defined as
@@
To couple cochains of adjacent dimensions, we introduce the Dirac-Bianconi operator bianconi2021topological
@@
In this work, we choose the orientation that allows for the homogeneous state, i.e., $\mathcal{D}\cdot\vec{1}=\vec{0}$, where 1=($\vec{1}=(1,...,1)^{\top}$ is a homogeneous vector of all ones, and $\vec{0}$ is the vector of all zeros. Note that this is not possible for every simplicial complex. For example, in the case of networks, only for Eulerian ones, i.e., those whose nodes have all even degree (see Giambagli2022diffusion for details) admit the homogeneous state. Hence, those are the kinds of networks that we are going to consider.
By applying the Dirac-Bianconi operator bianconi2021topological to the topological spinor, we can project the $0-$cochains onto the $1-$simplices and the $1-$coachains onto the $0-$simplices
@@
The new topological spinor $\vec{\psi}$ has a vector of $1-$cochain as a first component, which are the $0-$cochains projected onto the $1-$simplices, and a vector of $0-$cochains as a second component.
Note that, by taking the square of the Dirac-Bianconi operator, one obtains the Hogde Laplacian, which is the operator governing higher-order diffusion bianconi2021topological . As we will not deal with diffusion in this work, we will not discuss any further the Hodge Laplacian, whose detailed description can be found in the following Refs. bianconi2021higher ; Lim2020 .
@@ III Dirac-Bianconi driven oscillators
Let us consider a network of $N_{n}$ nodes and $N_{l}$ links, whose $N_{n}\times N_{n}$ adjacency matrix is given by $A$, and let us define the $N_{n}\times N_{l}$ boundary operator $\bm{B}_{1}$ through the orientation that admits the homogeneous state. Let us consider the 00-cochain $\vec{u}$, column vector of dimension $N_{n}$, which denotes the topological signals in the space of the nodes, i.e., the 00-simplices, and the $1$-cochain $\vec{v}$, column vector of dimension $N_{l}$, denoting the topological signals in the space of the links, i.e., the $1$-simplices. Note that each cochain has dimension [^3] In principle, each 00-cochain $u_{i}$ could be a $d$-dimensional vector, meaning that $\vec{u}$, the vector encoding all the 00-cochains of the system, would have dimension $dN_{n}$. For instance, one could deal with $2$-dimensional 00-cochains and $1$-dimensional $1$-cochains, as in muolo2024three . $1$. Going back to our system, the state variables of the whole network are described by the $N_{n}+N_{l}$-dimensional topological spinor $\vec{w}=(\vec{u}^{\top},\vec{v}^{\top})^{\top}$. The dynamics of the topological spinor is given by the following equation
@@
where $\vec{F}=(\vec{f}^{\top},\vec{g}^{\top})^{\top}$, and $f_{i}$ (resp. $g_{j}$) acts element-wise on the 00-cochains $u_{i},(\bm{B}_{1}\vec{v})_{i}$, for $i=1,...,N_{n}$ (resp. $1$-cochains $v_{j},(\bm{B}_{1}^{\top}\vec{u})_{j}$, for $j=1,...,N_{l}$).
@@
Looking at the above equations, one can observe that the dynamical variables, i.e., the 00- and $1$-cochains, interact with the variables on other simplices solely due to the Dirac-Bianconi operator. Otherwise, one would have $N_{n}$ decoupled systems on the nodes and $N_{l}$ decoupled systems on the links. This scheme is depicted in Fig. 1 for $N_{n}=N_{l}=3$. Note that this is different from the setting considered, for instance, in Giambagli2022diffusion , where the presence of a diffusion coupling through the Hodge Laplacian was ensuring that the variables on the nodes were coupled to each other and the variables on the links as well, while the Dirac operator was responsible for the inter-dimensional coupling. Here, without the Dirac-Bianconi operator, there would be no coupling at all.
@@
Figure 1: Visualization of a Dirac-Bianconi driven oscillator: the 00-cochains $u_{i}$ and $1$-cochains $v_{i}$, grouped in the topological spinor $\vec{w}$, are isolated and do not feel each other’s presence until the Dirac-Bianconi operator $\mathcal{D}$ couples them together, giving rise to a network whose dynamics is periodic. The oscillations are the results or the interactions between adjacent 00- and $1$-cochains through the boundary and co-boundary operators $\bm{B}_{1}$ and $\bm{B}_{1}^{\top}$, i.e., the Dirac-Bianconi operator.
@@
Let us now consider a solution of system (6), which we denote with $\vec{\mathcal{W}}$. We will call system (6) a Dirac-Bianconi driven oscillator when at least one of its solutions exhibit a stable periodic behavior, namely, when there exists a period $T$ such that $\vec{\mathcal{W}}(t+T)=\vec{\mathcal{W}}(t)$ for all time $t$. It is important to note that such periodic solution, which is driven by the Dirac-Bianconi coupling, is also a solution of the whole system. Each isolated unit, 00- or $1$-cochain, does not exhibit a periodic behavior. In other words, we defined as a Dirac-Bianconi driven oscillator a system of decoupled inter-dimensional units which exhibit a stable periodic behavior when coupled via the Dirac-Bianconi operator. For the sake of simplicity, we have considered (and we will consider) the case of 00- and $1$-cochains (i.e., nodes and links), but our definition can be extended to higher dimensional systems, i.e., $n$- and $n+1$-cochains. We chose the name Dirac-Bianconi driven oscillators to distinguish them from self-sustained oscillatory cochains, which are sometimes called topological oscillators. A topological oscillator is a self-sustained oscillator, meaning that it oscillates even without coupling, an example of such being the Topological Kuramoto millan2020explosive , where each variable on the links (or any other simplex) is an oscillator. However, in our case, the oscillatory dynamics is induced by the Dirac-Bianconi operator, hence the name Dirac-Bianconi driven oscillator.
@@
As anticipated in the Introduction, it was already observed that a pure Dirac-Bianconi coupling can yield patterns Giambagli2022diffusion , which can also be time-periodic muolo2024three . Indeed, the Dirac-induced patterns described in the latter Reference are an example of a Dirac-Bianconi driven oscillator, as it falls within the definition hereby given. In this work, we consider a more general setting with respect to the study mentioned above [^4] Ref. muolo2024three falls within the framework of Turing patterns Turing , where a homogeneous equilibrium is needed to develop the theory. When the setting is extended to networks and higher-order structures muolo2024turing , such conditions translates into considering identical oscillators and, if the state variables are cochains, additional conditions on the simplicial complex’ structure and the basis in which the boundary operators are defined (see Ref. carletti2023global for a detailed discussion)..

### v2

@@ I Introduction
In complex systems, the dynamics is shaped by the interactions among the elementary units and such interactions have been mostly assumed to be pairwise (two-body). However, in recent years, increasing evidence emerged towards the fundamental role that higher-order interactions play in describing real-world systems Battiston et al. (2020); Bianconi (2021a); Bick et al. (2023); Boccaletti et al. (2023). This triggered the interest of scholars in nonlinear dynamics, particularly due to the interplay between dynamics and higher-order topology Battiston et al. (2021); Muolo et al. (2024a); Millán et al. (2025). In fact, it has been shown that higher-order interactions enrich the dynamics of the systems, with applications regarding synchronization Tanaka and Aoyagi (2011); Skardal and Arenas (2020); Millán et al. (2020); Lucas et al. (2020); Wang et al. (2021); Skardal et al. (2021); Gambuzza et al. (2021); Gallo et al. (2022), random walks Schaub et al. (2020); Carletti et al. (2020a), pattern formation Carletti et al. (2020b); Muolo et al. (2023), chimera states Kundu and Ghosh (2022); Muolo et al. (2024b), opinion dynamics Iacopini et al. (2019); DeVille (2021); Ziegler et al. (2022); Schawe and Hernández (2022), and control De Lellis et al. (2022); Della Rossa et al. (2023); Muolo et al. (2025), to name a few. In all of the above works, the authors showed that certain dynamics are only made possible by the higher-order interaction topology, namely, hypergraphs and simplicial complexes, which are extensions of networks. This is particularly interesting, as the complexity of the system lies not in the model, but rather in the interactions between the elementary units, and even very simple settings yield rich dynamics when higher-order interactions are present Kuehn and Bick (2021); León et al. (2024).
@@
In the study of dynamics on networks, the classical approach relies on node-based dynamics, where to each node is associated a variable, e.g., the phase of an oscillator, or the state of a neuron. Higher-order interactions in the sense of many-body (group) interactions persist in such node-based paradigm: in fact, the dynamical variables remain anchored to the nodes, while what change is the interactions between them, mediated by hyperedges of higher orders, i.e., group interactions. However, there are systems which do not fit well in this paradigm, and some examples come from collaboration networks Patania et al. (2017), brain networks Faskowitz et al. (2022), climate science Singh et al. (2020), and transportation networks Katifori et al. (2010). To describe and analyze such systems, scholars have developed a framework that goes beyond node-based dynamics, using tools from algebraic topology and exterior calculus Bianconi (2021a), which allow us to consider the dynamics of topological signals, i.e., state variables defined not only on the nodes, but also on links, triangles, and higher-order simplices. The structure on which the dynamics is defined is called simplicial complex, which is a particular case of hypergraph with a closure relation, i.e., each $(d+1)$-dimensional simplices needs to be bounded by all $d$-simplices, and so on. As a simple example, a $2$-simplex (i.e., a triangle), needs to be bounded by all $1$-simplices (i.e., the links), which have to be bounded by all the 0-simplices (i.e., the nodes). In this notation, a network can be considered as a $1$-simplicial complex, i.e., a simplicial complex made only of $1$- and 0-simplices. Moreover, though the Dirac-Bianconi operator Bianconi (2021b), variables on different simplices of $1$ dimension up or down can be coupled. In this new framework, several classical nonlinear dynamics on networks model, such as the Kuramoto model Kuramoto (1975), Benjamin-Feir instability Nakao (2014), the Master Stability Function Fujisaka and Yamada (1983); Pecora and Carroll (1998), and Turing pattern formation Nakao and Mikhailov (2010); Muolo et al. (2024a), have been extended for topological signals on simplicial complexes, namely, the Topological Kuramoto model Millán et al. (2020); Calmon et al. (2022); Ghorbanchian et al. (2021); Li et al. (2023); Arnaudon et al. (2022); Nurisso et al. (2024); Wang et al. (2025), Global Topological Synchronization Carletti et al. (2023); Wang et al. (2024); Carletti et al. (2025), and Topological Turing patterns Giambagli et al. (2022); Muolo et al. (2024c). The study of patterns and oscillations of topological signals in higher-order simplices is particularly interesting for the study of edge-signals in the brain Faskowitz et al. (2022).
@@
In this work, we introduce a setting that is complementary to the above works. Before that, let us take a step back to the node-based framework. When studying synchronization dynamics on networks, it is common to view each node as a self-sustained oscillator, i.e., with a stable limit cycle, and then consider the synchronization between them Pikovsky et al. (2001). When considering such systems, there is a powerful technique, called phase reduction, which allows us to reduce each high dimensional oscillator to a phase variable Nakao (2016); Pietras and Daffertshofer (2019). In general, the application of phase reduction yields simple phase models that are easier to understand and that have been widely studied. In fact, the Kuramoto model is derived through phase reduction Kuramoto (1984) and provides a simple and general mechanism for the emergence of synchronization Kuramoto (1984); Acebrón et al. (2005); Rodrigues et al. (2016). In recent years, extensions of this method have been considered, such as corrections via an expansion of the (small) coupling strength León and Pazó (2019, 2022); Erik T. K. Mau (2023); Bick et al. (2023), the dynamics of the amplitude variables under non-weak coupling Kurebayashi et al. (2013); Wilson and Moehlis (2016); Shirasaka et al. (2017), or mean field variables León and Pazó (2020), which allowed us to overcome some of the limitations and extend its applicability. All the above references consider the nodes as oscillators. However, periodic dynamics can emerge also in more complex settings. For example, most neurons are not oscillatory, but oscillatory patterns are observed at larger scale Wilson and Cowan (1972); Brunel and Hakim (1999). This can be modeled by a network of interacting excitable units, whose isolated dynamics are not oscillatory, but exhibit self-sustained oscillations as a result of the coupling Kawamura et al. (2011); Pazó and Montbrió (2016); Devalle et al. (2017); Nakao et al. (2018); Politi et al. (2018); Di Volo and Torcini (2018); Nakao et al. (2021). In particular, authors of Ref. Nakao et al. (2018) derived a phase description of oscillatory networks and studied their synchronization through phase reduction theory.
@@
Inspired by the latter work, we further extend this framework by considering networks where the dynamical variables lie on the nodes and the links, and are solely coupled via the Dirac-Bianconi operator. When such systems exhibit self-sustained oscillations, we call them Dirac-Bianconi driven oscillators, because their periodic behavior is induced by the Dirac-Bianconi operator. We then give a full description of such structures and study their synchronization through phase reduction theory.
@@
The paper is organized as follows. In the next Section, we will give the theoretical background on topological signals and the Dirac-Bianconi operator. In Section III, we first introduce and characterize a Dirac-Bianconi driven oscillator, and then study numerically the synchronization of two coupled oscillators. Finally, in Sec. IV, we give a phase description of such systems that explains the numerical results.
@@ II Theory of topological signals and the Dirac-Bianconi operator
In this Section we present a basic introduction to simplicial complexes and topological signals, limiting ourselves to define what is needed to characterize Dirac-Bianconi driven oscillators. For a more detailed discussion that includes other important concepts, the interested reader may consult more specialized references Bianconi (2021a); Millán et al. (2025); Lek-Heng (2020); Schaub et al. (2021).
A $d$-dimensional simplex is a set of $d+1$ nodes that models a higher-order interaction among $(d+1)$ entities or represents a discrete object of dimension $n$. A 0-simplex is a node, a $1$-simplex is a link, a $2$-simplex is a triangle, and so on. The faces of a $d$-simplex are all its (d1)(d-1)-dimensional subsimplices. For instance, the faces of a triangle (2-simplex) are its three links (1-simplices), and each link has two node (0-simplex) faces. Except for the 0-simplices, each simplex has an arbitrary orientation, i.e., each simplex is labeled by an ordered list of nodes. For example, the (symmetric) link connecting nodes $i$ and $j$ can be labeled as $[i,j]$; then $[j,i]=[i,j]$. Linear combinations of $d$-simplices are also called $d$-chains. A simplicial complex is built by attaching simplices by their faces. Note that, in a simplicial complex, the existence of a $d$-simplex implies the existence of all its faces. Given $N_{k}$ $k$-simplices and $N_{k-1}$ (k1)(k-1)-simplices, the boundary operator $\mathbf{B}_{k}$ is a $N_{k-1}\times N_{k}$ matrix defined by:
[…]
Note that the boundary operator $\mathbf{B}_{k}$ changes when the orientation is changed, but the structure of the simplicial complex does not change; one can understand this by thinking of Cartesian coordinates: the axis can be changed, but this does not change the structure of the space. Note also that, in the case of network, i.e., consisting only of 0-simplices and $1$-simplices, $\mathbf{B}_{1}$ is the well-known incidence matrix Newman (2010).
We now define topological signals (also called cochains) as functions defined on simplices (chains). A $k$-dimensional signal is a function assigning a real value to each $k$-simplex $\sigma^{i}_{k}$. The values can be arranged in a vector [^1] Note that the dimension of the topological signal and the dimension of the simplex are not related. As one can define an $n$-dimensional vector on the nodes (0-simplices), the same can be done on any $k$-simplex. $\vec{x}\in\mathbb{R}^{N_{k}}$, where each entry is $x_{i}=x(\sigma^{i}_{k})$. These signals must be orientation-sensitive, i.e., $x(\sigma_{k})=-x(-\sigma_{k})$, to ensure the invariance under the change of the arbitrary orientation Carletti et al. (2023). Note that $d$-cochains are the discrete equivalent of differential $d$-forms Lek-Heng (2020); Grady and Polimeni (2010). Hence, 0-cochains naturally model scalar fields, $1$-cochains vector fields, e.g., the gradient of a scalar, $2$-cochains bivector fields, e.g., the curl of a vector Dorst et al. (2007).
@@
The boundary operator (1) connects topological signals (cochains) of consecutive dimensions: $\mathbf{B}_{k}$ on a $k$-cochain $\vec{x}$, it gives a (k1)(k-1)-cochain $\vec{z}=\mathbf{B}_{k}\vec{x}$, while acting with the transpose gives a $(k+1)$-cochain $\vec{g}=\mathbf{B}_{k+1}^{\top}\vec{x}$. The transpose of the boundary is known as the coboundary operator. One cannot connect topological signals of dimensions $k$ and $k\pm 2$ because the (co)boundary of a (co)boundary is zero Bianconi (2021a).
For sake of simplicity, let us consider the case of a $1$-simplicial complex, i.e., a network, hence, only $0$-simplices (nodes) and 11-simplices (links) are involved, with their respective topological signals [^2] Note that the terms cochains, which we will mainly use from now on, and topological signals are equivalent., namely, $0$-cochains $\vec{u}$ (state variables on the nodes) and $1$-cochains $\vec{v}$ (state variables on the links). Note that the same formalism can be used to consider any dimension, as long as they are consecutive, given that Dirac-Bianconi operator connects topological signals of consecutive dimension. The state of the system is described by the topological spinor $\vec{w}$, defined as
@@
To couple cochains of adjacent dimensions, we introduce the Dirac-Bianconi operator Bianconi (2021b)
@@
In this work, we choose the orientation that allows for the homogeneous state, i.e., $\mathcal{D}\cdot\vec{1}=\vec{0}$, where $\vec{1}=(1,...,1)^{\top}$ is a homogeneous vector of all ones, and $\vec{0}$ is the vector of all zeros. Note that this is not possible for every simplicial complex. For example, in the case of networks, only for Eulerian ones, i.e., those whose nodes have all even degree (see Giambagli et al. (2022) for details) admit the homogeneous state. Hence, those are the kinds of networks that we are going to consider.
By applying the Dirac-Bianconi operator Bianconi (2021b) to the topological spinor, we can project the $0$-cochains onto the $1$-simplices and the 11-coachains onto the $0$-simplices
@@
The new topological spinor $\vec{\psi}$ has a vector of 11-cochain as a first component, which are the $0$-cochains projected onto the $1$-simplices, and a vector of $0$-cochains as a second component.
Note that, by taking the square of the Dirac-Bianconi operator, one obtains the Hogde Laplacian, which is the operator governing higher-order diffusion Bianconi (2021b). As we will not deal with diffusion in this work, we will not discuss any further the Hodge Laplacian, whose detailed description can be found in the following Refs. Bianconi (2021a); Lek-Heng (2020).
@@ III Dirac-Bianconi driven oscillators
Let us consider a network of $N_{n}$ nodes and $N_{l}$ links, whose $N_{n}\times N_{n}$ adjacency matrix is given by $A$, and let us define the $N_{n}\times N_{l}$ boundary operator $\bm{B}_{1}$ through the orientation that admits the homogeneous state. Let us consider the 0-cochain $\vec{u}$, column vector of dimension $N_{n}$, which denotes the topological signals in the space of the nodes, i.e., the 0-simplices, and the $1$-cochain $\vec{v}$, column vector of dimension $N_{l}$, denoting the topological signals in the space of the links, i.e., the $1$-simplices. Note that each cochain has dimension [^3] In principle, each 0-cochain $u_{i}$ could be a $d$-dimensional vector, meaning that $\vec{u}$, the vector encoding all the 0-cochains of the system, would have dimension $dN_{n}$. For instance, one could deal with $2$-dimensional 0-cochains and $1$-dimensional 11-cochains, as in Muolo et al. (2024c). $1$. Going back to our system, the state variables of the whole network are described by the $N_{n}+N_{l}$-dimensional topological spinor $\vec{w}=(\vec{u}^{\top},\vec{v}^{\top})^{\top}$. The dynamics of the topological spinor is given by the following equation
@@
where $\vec{F}=(\vec{f}^{\top},\vec{g}^{\top})^{\top}$, and $f_{i}$ (resp. $g_{j}$) acts element-wise on the 0-cochains $u_{i},(\bm{B}_{1}\vec{v})_{i}$, for $i=1,...,N_{n}$ (resp. $1$-cochains $v_{j},(\bm{B}_{1}^{\top}\vec{u})_{j}$, for $j=1,...,N_{l}$).
@@
Looking at the above equations, one can observe that the dynamical variables, i.e., the 0- and $1$-cochains, interact with the variables on other simplices solely due to the Dirac-Bianconi operator. Otherwise, one would have $N_{n}$ decoupled systems on the nodes and $N_{l}$ decoupled systems on the links. This scheme is depicted in Fig. 1 for $N_{n}=N_{l}=3$. Note that this is different from the setting considered, for instance, in Giambagli et al. (2022), where the presence of a diffusion coupling through the Hodge Laplacian was ensuring that the variables on the nodes were coupled to each other and the variables on the links as well, while the Dirac operator was responsible for the inter-dimensional coupling. Here, without the Dirac-Bianconi operator, there would be no coupling at all.
@@
Figure 1: Visualization of a Dirac-Bianconi driven oscillator: the 0-cochains $u_{i}$ and $1$-cochains $v_{i}$, grouped in the topological spinor $\vec{w}$, are isolated and do not feel each other’s presence until the Dirac-Bianconi operator $\mathcal{D}$ couples them together, giving rise to a network whose dynamics is periodic. The oscillations are the results or the interactions between adjacent 0- and $1$-cochains through the boundary and co-boundary operators $\bm{B}_{1}$ and $\bm{B}_{1}^{\top}$, i.e., the Dirac-Bianconi operator.
@@
Let us now consider a solution of system (6), which we denote with $\vec{\mathcal{W}}$. We will call system (6) a Dirac-Bianconi driven oscillator when at least one of its solutions exhibit a stable periodic behavior, namely, when there exists a period $T$ such that $\vec{\mathcal{W}}(t+T)=\vec{\mathcal{W}}(t)$ for all time $t$. It is important to note that such periodic solution, which is driven by the Dirac-Bianconi coupling, is also a solution of the whole system. Each isolated unit, 0- or $1$-cochain, does not exhibit a periodic behavior. In other words, we defined as a Dirac-Bianconi driven oscillator a system of decoupled inter-dimensional units which exhibit a stable periodic behavior when coupled via the Dirac-Bianconi operator. For the sake of simplicity, we have considered (and we will consider) the case of 0- and $1$-cochains (i.e., nodes and links), but our definition can be extended to higher dimensional systems, i.e., $n$- and $n+1$-cochains. We chose the name Dirac-Bianconi driven oscillators to distinguish them from self-sustained oscillatory cochains, which are sometimes called topological oscillators. A topological oscillator is a self-sustained oscillator, meaning that it oscillates even without coupling, an example of such being the Topological Kuramoto Millán et al. (2020), where each variable on the links (or any other simplex) is an oscillator. However, in our case, the oscillatory dynamics is induced by the Dirac-Bianconi operator, hence the name Dirac-Bianconi driven oscillator.
@@
As anticipated in the Introduction, it was already observed that a pure Dirac-Bianconi coupling can yield patterns Giambagli et al. (2022), which can also be time-periodic Muolo et al. (2024c). Indeed, the Dirac-induced patterns described in the latter Reference are an example of a Dirac-Bianconi driven oscillator, as it falls within the definition hereby given [^4] Ref. Muolo et al. (2024c) falls within the framework of Turing patterns Turing (1952), where a homogeneous equilibrium is needed to develop the theory. When the setting is extended to networks and higher-order structures Muolo et al. (2024a), such conditions translates into considering identical oscillators and, if the state variables are cochains, additional conditions on the simplicial complex’ structure and the basis in which the boundary operators are defined (see Ref. Carletti et al. (2023) for a detailed discussion).. There, a $2$-dimensional system lays in the nodes (i.e., two 0-cochains on the 0-chains) in a stationary state and oscillations were induced by the presence of third species on the links (i.e., a $1$-cochain on the $1$-chains). Note that, when more than one $d$-cochains lie on their respective $d$-chains, oscillations are possible for some given vector fields. In principle, one could further extend the framework of Ref. Muolo et al. (2024c) by considering a $2$-dimensional system on the nodes and another $2$-dimensional system on the links, both in a stationary state, and induce oscillations via the Dirac-Bianconi coupling. However, although the Dirac-Bianconi coupling is what triggers the emergence of oscillations, periodic behaviors would still be possible for appropriate values of the parameters. In this work, instead, we consider a setting in which any oscillatory dynamics would be impossible, if not for the Dirac-Bianconi coupling.